# EPOCHS IV: SED Modelling Assumptions and their impact on the Stellar Mass Function at using PEARLS and public JWST observations

### v1

@@ 6.4 Impact of Modified IMF
Our implementation of a modified top-heavy IMF in Prospector has been shown to reduce masses by up to 0.5 dex for galaxies at 12 (HOT 60K) when compared to a standard Kroupa (2001) IMF, as shown in Figure 5. The decrease in mass seen is also found to be dependent on the star formation model, with a larger decrease in stellar mass found when assuming a parametric “delayed” SFH model (0.46 dex) than when assuming a non-parametric “continuity bursty” SFH model (0.35 dex). Comparison of the best-fitting models has also shown that the is almost unaffected by the change in IMF, suggesting the modified IMF models can match the observed photometry of the galaxies as well as the original data. We see similar results, with smaller decreases in stellar mass (0.3 dex), for the modified IMF model used at 8 (HOT 45K).
There are a number of comparable studies which also look at the impact of a top-heavy IMF on the masses of high redshift galaxies. The most direct comparison can be made to the results of Steinhardt et al. (2023) as we have used the same top-heavy IMF model with a different SED fitting tool. Their results are based on some of the earliest JWST observations and they only fit a small number of galaxies. Direct comparison of their results to the default EAZY-py fsps templates is also impacted by other differences between the two template sets, including the re-parametrization of the IMF from Chabrier et al. (2000) to Kroupa (2001) and different assumptions for SFH and nebular emission. Our implementation changes only the IMF, with everything else modelled in the same way. They observe decreases of dex in stellar mass, whereas our implementation sees smaller decreases, typically from dex. The reason for this discrepancy is unclear, but it may be possible that the other differences between the standard EAZY-py templates and their models have a larger effect on the stellar mass than expected, or that the change in IMF in our modelling resulted in different star formation histories that acted to somewhat decrease the impact of the modified IMF.
Woodrum et al. (2023) also look at other top-heavy modifications to commonly used IMF parametrizations using Prospector. They look at a modification to the Chabrier et al. (2000) IMF, rather than the Kroupa (2001) we use, but they find comparable reductions in stellar mass (between 0.38 and 0.5 dex). They also see no change in the goodness of fit when using a modified IMF.
Whilst we have shown that a modified top-heavy IMF can decrease the stellar masses of high- galaxies, with no change in the simulated photometry, our analysis of the CDM limits on the growth of stellar mass in § 6.1 does not require a non-standard IMF in order to be compatible with CDM. In Figure 1 we have shown that the stellar masses can vary significantly with other SED fitting assumptions (dust law, assumed SFH), before any IMF change is considered. For the galaxies with high stellar masses across all the models we consider, it is not possible to distinguish between a modified IMF or simply a high star formation efficiency based on photometry alone.
@@ 6.5 Comparing the measured GSMF with other observations and theory/simulations
In Figure 8, we compare our GSMF estimates to a wide-range of observational and theoretical/simulation based predictions of the GSMF. Here we briefly discuss our GSMF estimates for each of the redshift bins. In order to make direct comparisons where necessary all results have been converted to use a Kroupa (2001) IMF.
The overall evolution of the derived Schechter parameters and a comparison to the results derived by other studies can be seen in Figure 10. Whilst the Schechter parameters are highly covariant, and our results typically have large uncertainties, we observe an evolution in and compared, with both parameters decreasing compared to the results at of Caputi et al. (2015), Duncan et al. (2014) and Grazian et al. (2015). We see little evolution of M within the large range of uncertainties.
@@ 6.5.1 Redshift z=7 GSMF
We derive a mass function at primarily as a proof of concept of our method. We do not take advantage of galaxy lensing in this work, so any reasonable mass completeness limit is higher than previous studies, nor do we have the area of wide-field studies like Weaver et al. (2022) in order to detect rare, bright and high mass galaxies. However with JWST we have seen a surprising excess of UV-faint LRD like objects with high inferred stellar masses, as discussed in § 6.3. The majority of these sources were previously undetected with HST due to relatively weak Lyman-breaks, and so do not appear in pre-JWST stellar mass estimates. Their inclusion in our GSMF has resulted in an excess at the high mass end of the GSMF when compared to other observational studies, and consequently a higher and poorly constrained estimate of M, as we see little evidence for any exponential turnover. The highest mass GSMF data-points of Weaver et al. (2022) fall within our 1 uncertainty region, but our results are significantly above the measurements of Stefanon et al. (2021). At the low-mass end we fall below the results of Kikuchihara et al. (2020) and Navarro-Carrera et al. (2023), but agree within the uncertainties of Furtak et al. (2021) and Bhatawdekar et al. (2019). Stefanon et al. (2021); Kikuchihara et al. (2020); Bhatawdekar et al. (2019) and Furtak et al. (2021) are all based on HST+Spitzer observations of the Hubble Frontier Fields, and incorporate lensing, which means they probe the low mass end of the GSMF more accurately then this study. Our low mass slope is in good agreement with Furtak et al. (2021), but steeper than the results Stefanon et al. (2021) and Kikuchihara et al. (2020). At the time of writing, Navarro-Carrera et al. (2023), Gottumukkala et al. (2023) and Wang et al. (2024) are the only other studies to incorporate JWST observations into their GSMF estimates. We see reasonable agreement with the results of Navarro-Carrera et al. (2023) as our data is within their GSMF uncertainties. This work relies almost entirely on JWST observations, whereas they combine deep JWST observations of small volumes (20 arcmin) with HST and ground-based catalogues. This ground-based data allows them to find more rare, high-mass galaxies than our study, but at lower masses the small volumes probed with their JWST data are potentially vulnerable to cosmic variance. Reliance on ground-based and HST data also limits the maximum redshift they can probe to . Wang et al. (2024) uses PRIMER observations with NIRCam and MIRI to measure the GSMF, notably finding that the use of MIRI observations systematically reduces stellar masses measured with SED fitting. We see good agreement in the measured GSMF within the uncertainties of both studies, despite this work not incorporating MIRI data or correcting for any systematic offset in mass arising from the lack of restframe m observations. Our GSMF does extend to higher stellar mass than the result of Wang et al. (2024), resulting in a higher value for , as can be seen in Figure 10.
@@
Gottumukkala et al. (2023) examine the contribution of high-mass, dusty galaxies at to the GSMF using data from the CEERS survey. Given that our GSMF probes a wider galaxy population, we do not expect to see overlap at all stellar masses. We see good overlap at the highest stellar masses M, where our SMF estimate is dominated by dusty LRD galaxies (as discussed in § 6.3).
@@ 6.5.2 Redshift z=8 GSMF
Our fiducial GSMF estimate at shows reasonable agreement with most predictions. As we do not bootstrap in redshift when constructing the GSMF, we do not account for galaxies scattering between redshift bins, and for example a galaxy found to be at with Bagpipes would contribute only to the GSMF, even in a significant fraction of the redshift PDF is above . This does not affect the majority of galaxies within our sample, but it does explain some of the discrepancy between our results and the implied results of Labbé et al. (2023), shown in purple in Figure 8 assuming 100% completeness. The LRD galaxies of Labbé et al. (2023) which we also select are all found to be at , meaning they do not contribute at all to our estimate of the GSMF. The best-fitting redshift for these objects in some cases is quite close to this boundary however, meaning that these objects could theoretically contribute to the GSMF instead, which would boost the high-mass end significantly. Our GSMF is also lower than the results of Kikuchihara et al. (2020), which incorporates strong gravitational lensing in order to probe to lower stellar mass. Our GSMF agrees with the results of Song et al. (2016); Bhatawdekar et al. (2019); Stefanon et al. (2021), and appears to validate the majority of pre-JWST GSMF estimates. We can also draw a comparison to the results of Wang et al. (2024), whose GSMF results at are systematically above our results, but within the derived uncertainties of both studies.
@@ 6.5.3 Redshift z=9 GSMF
Our GSMF estimate at is below the results of Bhatawdekar et al. (2019) and Kikuchihara et al. (2020), but within the uncertainties of Stefanon et al. (2021). We are below the implied result of Labbé et al. (2023), which is derived from two galaxies in their sample in this redshift bin, but assuming 100% completeness. We include one of these galaxies in our GSMF at this redshift, as the other does not meet our selection criteria, as we do not detect the Lyman break at 5. We additionally include one other candidate from their sample in this GSMF, as our fiducial Bagpipes photo- places it within this redshift bin, rather than the redshift bin based on their photo-. For both of their galaxies we do include in this redshift bin we find dex lower stellar masses, meaning they contribute to a lower stellar mass bin. Our reliance on the rest-frame UV to robustly detect sources is one limitation of this work, although the increased depth of JWST observations when compared to HST has reduced this in some fields. We investigated less-stringent constraints on the Lyman-break, but found that this dramatically increased rates of contamination within our sample. When compared to simulations, FLARES and Universe Machine are close to our GSMF estimate, but almost all the predictions are within our posterior region. Interestingly, in this redshift bin we are below the predictions of two recent JWST-era studies; Mauerhofer & Dayal (2023) and Li et al. (2023), both of which incorporate higher star formation efficiencies than typical models.
@@
In our highest redshift bin, 11.5 which covers only 80 Myr, there are no published observationally-derived results and few theoretical or simulation-based GSMF comparisons at this redshift. Pre-JWST estimates of the GSMF were not possible at this redshift, and even with JWST our GSMF estimate is also uncertain due to the difficulty in accurate stellar mass estimates as well as the possible contribution of contaminants. At the longest wavelength NIRCam filter falls within the rest-frame UV, which is dominated by young stars, leading to highly uncertain star formation histories and stellar masses. As we explored briefly in § 4.5, the more likely possibility of a top-heavy IMF or exotic stellar populations in these early galaxies further increases the systematic uncertainties in the stellar mass estimates. An example of three galaxies within this redshift bin is shown in the lower plot of Figure 3 and in Figure 4, and the range of stellar mass estimates () for different Bagpipes and Prospector with very little difference in the fitted rest-UV spectra shows the difficulty in estimating stellar mass at these redshift. Previous studies at lower redshift with HST and Spitzer have found that stellar masses estimated by HST alone, with no measurement of the rest optical emission, typically underestimate stellar masses by 0.62 dex, compared to measurements including HST and Spitzer NIR observations (Furtak et al., 2021). At this redshift range, our JWST NIRCam observations are probing comparable rest-frame UV wavelengths to HST observations at , and it possible that our stellar masses are also underestimated unless there is a significant change in stellar populations or IMF. The possibility of more stochastic star formation histories at this redshift compared to may also lead to outshining, which further increases the stellar mass discrepancy (Narayanan et al., 2023).
@@
We note that several galaxies in this bin have been excluded from the GSMF in this case due to our requirement that the contamination is less than 50%. The inclusion of these possible contaminants would result in an 0.3 dex increase in the lowest mass bin. The results of this are shown in Appendix C. Whilst we do not attempt to fit the GSMF, we can make approximate comparisons to the few available predictions. We see the closest agreement with DELPHI (Mauerhofer & Dayal, 2023) and are within 1.5 of FLARES (Lovell et al., 2021; Wilkins et al., 2023b) in the higher mass bin, but have an excess of M galaxies when comparing to FLARES and SC SAM GUREFT (Yung et al., 2023). Our results are significantly above the predictions of BLUETIDES (Feng et al., 2016; Wilkins et al., 2017) at all stellar masses. Our fiducial GSMF prediction is slightly below the prediction of Li et al. (2023) with maximum star formation efficiency = 0.2, and significantly below the upper limit of = 1.
@@ 6.5.6 Alternative GSMF Estimates
Resolved SED fitting from Giménez-Arteaga et al. (2023) has found that the higher stellar masses inferred by the non-parametric SFHs can better account for the out-shining of older stellar populations in bright, actively star forming galaxies. How widespread the issue of outshining is among our sample is unknown, but if it is common then the masses and resulting GSMF of the “continuity bursty” SFH model may more accurately recover the true stellar populations of these early galaxies.
@@ 6.6 Stellar Mass Density Evolution in the Early Universe
The growth of stellar mass density in the early Universe is highly uncertain. Some observational studies (e.g. Oesch et al., 2014; Stefanon et al., 2021; Willott et al., 2023a) have found a sharp decline in stellar mass density at 8, whereas others see a flatter evolution (Kikuchihara et al., 2020; Bhatawdekar et al., 2019). On the theoretical side, Oesch et al. (2018) uses dark matter halo evolutionary models to predict a deviation from the constant star formation efficiency (CSFE) model of Madau & Haardt (2015), which follows a significantly steeper slope at 7.
@@
The range of stellar mass densities possible with our alternative GSMF estimates at is mostly above the 1 range of Adams et al. (2023). A discrepancy between the integrated star formation rate density and stellar mass density measured for the same sample could hint at a different IMF, since the assumed return fraction is strongly dependent on the chosen IMF, and the SMF and UVLF probe different stellar populations with different characteristic stellar mass. However there are a number of other possible issues with the conversion of the UVLF into a SMD estimate; the conversion of UV flux to SFR (), is often assumed to be a constant factor but actually dependent on the age and metallicity Madau & Dickinson (2014), as well as the other assumptions used to calculate the return fraction (closed-box model, constant IMF & metal yield and instantaneous recycling of metals) which may not be valid at high-redshift.

### v3

@@ 6.3 Impact of Modified IMF
Our modified top-heavy IMF implementation in Prospector reduces masses by up to 0.5 dex for galaxies at 12 (HOT 60K) compared to the standard Kroupa (2001) IMF, as shown in Figure 4 . The mass decrease depends on the star formation model, with a larger reduction seen in the parametric ”delayed” SFH model (0.46 dex) compared to the non-parametric ”continuity bursty” SFH model (0.35 dex). The best-fitting models show the is almost unaffected by the IMF change, indicating the modified IMF models match the observed photometry as well as the original data. For the modified IMF model used at 8 (HOT 45K), we see smaller mass decreases (0.3 dex).
Comparable studies also examine the impact of a top-heavy IMF on high-redshift galaxy masses. Steinhardt et al. (2023), using the same top-heavy IMF model with a different SED fitting tool, observed decreases of dex in stellar mass, while our implementation shows smaller decreases of dex. This discrepancy may be due to differences between the standard EAZY-py templates and their models, or variations in star formation histories resulting from the IMF change. Woodrum et al. (2023) also studied top-heavy IMF modifications using Prospector, focusing on a modification to the Chabrier et al. (2000) IMF rather than Kroupa (2001). They found similar reductions in stellar mass (0.38 to 0.5 dex) with no change in the goodness of fit.
Our analysis shows that a modified top-heavy IMF can decrease the stellar masses of high- galaxies without altering the simulated photometry. However, our examination of the CDM limits on stellar mass growth (§ 6.1 ) does not require a non-standard IMF for compatibility with CDM. As shown in Figure 1, stellar masses can vary significantly with other SED fitting assumptions (dust law, assumed SFH) before considering any IMF changes. For galaxies with high stellar masses across all models, it is challenging to distinguish between a modified IMF and high star formation efficiency based on photometry alone.
@@ 6.4 Comparing the measured GSMF with other observations and theory/simulations
In Figure 7, we compare our GSMF estimates to a wide-range of observational and theoretical/simulation based predictions of the GSMF. Here we briefly discuss our GSMF estimates for each of the redshift bins. In order to make direct comparisons where necessary all results have been converted to use a Kroupa (2001) IMF. The overall evolution of the derived Schechter parameters and a comparison to the results derived by other studies can be seen in Figure 9. Whilst the Schechter parameters are highly covariant, and our results typically have large uncertainties, we observe an evolution in and compared, with both parameters decreasing compared to the results at of Caputi et al. (2015), Duncan et al. (2014) and Grazian et al. (2015). We see little evolution of M⋆ within the large range of uncertainties.
@@ 6.4.1 Redshift z=7 GSMF
We derive a mass function at primarily as a proof of concept of our method. We do not take advantage of galaxy lensing in this work, so any reasonable mass completeness limit is higher than previous studies, nor do we have the area of wide-field studies like Weaver et al. (2022) in order to detect rare, bright and high mass galaxies. However with JWST we have seen a surprising excess of UV-faint LRD like objects with high inferred stellar masses, as discussed in § 6.2. The majority of these sources were previously undetected with HST due to relatively weak Lyman-breaks, and so do not appear in pre-JWST stellar mass estimates. Their inclusion in our GSMF has resulted in an excess at the high mass end of the GSMF when compared to other observational studies, and consequently a higher and poorly constrained estimate of M⋆, as we see little evidence for any exponential turnover. The highest mass GSMF data-points of Weaver et al. (2022) fall within our 1 uncertainty region, but our results are significantly above the measurements of Stefanon et al. (2021). At the low-mass end we fall below the results of Kikuchihara et al. (2020) and Navarro-Carrera et al. (2023), but agree within the uncertainties of Furtak et al. (2021) and Bhatawdekar et al. (2019). Stefanon et al. (2021); Kikuchihara et al. (2020); Bhatawdekar et al. (2019) and Furtak et al. (2021) are all based on HST+Spitzer observations of the Hubble Frontier Fields, and incorporate lensing, which means they probe the low mass end of the GSMF more accurately then this study. Our low mass slope is in good agreement with Furtak et al. (2021), but steeper than the results Stefanon et al. (2021) and Kikuchihara et al. (2020). At the time of writing, Navarro-Carrera et al. (2023), Gottumukkala et al. (2023), Wang et al. (2024) and Weibel et al. (2024) are the only other studies to incorporate JWST observations into their GSMF estimates. We see reasonable agreement with the results of Navarro-Carrera et al. (2023) as our data is within their GSMF uncertainties. This work relies almost entirely on JWST observations, whereas they combine deep JWST observations of small volumes (20 arcmin2) with HST and ground-based catalogues. This ground-based data allows them to find more rare, high-mass galaxies than our study, but at lower masses the small volumes probed with their JWST data are potentially vulnerable to cosmic variance. Reliance on ground-based and HST data also limits the maximum redshift they can probe to . Wang et al. (2024) uses PRIMER observations with NIRCam and MIRI to measure the GSMF, notably finding that the use of MIRI observations systematically reduces stellar masses measured with SED fitting. We see good agreement in the measured GSMF within the uncertainties of both studies, despite this work not incorporating MIRI data or correcting for any systematic offset in mass arising from the lack of restframe m observations. Our GSMF does extend to higher stellar mass than the result of Wang et al. (2024), resulting in a higher value for , as can be seen in Figure 9.
@@
Gottumukkala et al. (2023) examine the contribution of high-mass, dusty galaxies at to the GSMF using data from the CEERS survey. Given that our GSMF probes a wider galaxy population, we do not expect to see overlap at all stellar masses. We see good overlap at the highest stellar masses M, where our SMF estimate is dominated by dusty LRD galaxies (as discussed in § 6.2). We see good agreement with Weibel et al. (2024) at , who construct the GSMF at using data from JADES, CEERS and PRIMER.
@@ 6.4.2 Redshift z=8 GSMF
Our fiducial GSMF estimate at shows reasonable agreement with most predictions. As we do not bootstrap in redshift when constructing the GSMF, we do not account for galaxies scattering between redshift bins, and for example a galaxy found to be at with Bagpipes would contribute only to the GSMF, even in a significant fraction of the redshift PDF is above . This does not affect the majority of galaxies within our sample, but it does explain some of the discrepancy between our results and the implied results of Labbé et al. (2023), shown in purple in Figure 7 assuming 100% completeness. The LRD galaxies of Labbé et al. (2023) which we also select are all found to be at , meaning they do not contribute at all to our estimate of the GSMF. The best-fitting redshift for these objects in some cases is quite close to this boundary however, meaning that these objects could theoretically contribute to the GSMF instead, which would boost the high-mass end significantly. Our GSMF is also lower than the results of Kikuchihara et al. (2020), which incorporates strong gravitational lensing in order to probe to lower stellar mass. Our GSMF agrees with the results of Song et al. (2016); Bhatawdekar et al. (2019); Stefanon et al. (2021), and appears to validate the majority of pre-JWST GSMF estimates. We can also draw a comparison to the results of Wang et al. (2024), whose GSMF results at are systematically above our results, but within the derived uncertainties of both studies. We are also below the results of Weibel et al. (2024).
@@ 6.4.3 Redshift z=9 GSMF
Our GSMF estimate at is below the results of Bhatawdekar et al. (2019) and Kikuchihara et al. (2020), but within the uncertainties of Stefanon et al. (2021). We are below the implied result of Labbé et al. (2023), which is derived from two galaxies in their sample in this redshift bin, but assuming 100% completeness. We include one of these galaxies in our GSMF at this redshift, as the other does not meet our selection criteria, as we do not detect the Lyman break at 5. We additionally include one other candidate from their sample in this GSMF, as our fiducial Bagpipes photo- places it within this redshift bin, rather than the redshift bin based on their photo-. For both of their galaxies we do include in this redshift bin we find dex lower stellar masses, meaning they contribute to a lower stellar mass bin. Our reliance on the rest-frame UV to robustly detect sources is one limitation of this work, although the increased depth of JWST observations when compared to HST has reduced this in some fields. We investigated less-stringent constraints on the Lyman-break, but found that this dramatically increased rates of contamination within our sample. We are also below the results of Weibel et al. (2024), but within the uncertainties where they consider their GSMF reliable, at , as they suspect low-z contamination at higher inferred masses. When compared to simulations, FLARES and Universe Machine are close to our GSMF estimate, but almost all the predictions are within our posterior region. Interestingly, in this redshift bin we are below the predictions of two recent JWST-era studies; Mauerhofer & Dayal (2023) and Li et al. (2023), both of which incorporate higher star formation efficiencies than typical models.
@@
In our highest redshift bin, 11.5 which covers only 80 Myr, there are no published observationally-derived results and few theoretical or simulation-based GSMF comparisons at this redshift. Pre-JWST estimates of the GSMF were not possible at this redshift, and even with JWST our GSMF estimate is also uncertain due to the difficulty in accurate stellar mass estimates as well as the possible contribution of contaminants. At the longest wavelength NIRCam filter falls within the rest-frame UV, which is dominated by young stars, leading to highly uncertain star formation histories and stellar masses. As we explored briefly in § 4.4, the more likely possibility of a top-heavy IMF or exotic stellar populations in these early galaxies further increases the systematic uncertainties in the stellar mass estimates. An example of three galaxies within this redshift bin is shown in the lower plot of Figure 2 and in Figure 3, and the range of stellar mass estimates () for different Bagpipes and Prospector with very little difference in the fitted rest-UV spectra shows the difficulty in estimating stellar mass at these redshift. Previous studies at lower redshift with HST and Spitzer have found that stellar masses estimated by HST alone, with no measurement of the rest optical emission, typically underestimate stellar masses by 0.62 dex, compared to measurements including HST and Spitzer NIR observations (Furtak et al., 2021). At this redshift range, our JWST NIRCam observations are probing comparable rest-frame UV wavelengths to HST observations at , and it possible that our stellar masses are also underestimated unless there is a significant change in stellar populations or IMF. The possibility of more stochastic star formation histories at this redshift compared to may also lead to outshining, which further increases the stellar mass discrepancy (Narayanan et al., 2023). In order to test whether the reliance of the rest UV is impacting the stellar masses compared to lower redshifts, we refit all galaxies using Bagpipes in the bin without F410M or F444W, which gives equivalent rest-frame coverage to a galaxy. We see no systematic shift in the stellar masses, with the median offset being 0.07 dex and a standard deviation of 0.3 dex, when comparing to the stellar mass estimate including the longer wavelengths filters.
@@
We note that several galaxies in this bin have been excluded from the GSMF in this case due to our requirement that the contamination is less than 50%. The inclusion of these possible contaminants would result in an 0.3 dex increase in the lowest mass bin. The results of this are shown in Appendix D. Whilst we do not attempt to fit the GSMF, we can make approximate comparisons to the few available predictions. We see the closest agreement with DELPHI (Mauerhofer & Dayal, 2023) and are within 1.5 of FLARES (Lovell et al., 2021; Wilkins et al., 2023b) in the higher mass bin, but have an excess of M⊙ galaxies when comparing to FLARES and SC SAM GUREFT (Yung et al., 2023). Our results are significantly above the predictions of BLUETIDES (Feng et al., 2016; Wilkins et al., 2017) at all stellar masses. Our fiducial GSMF prediction is slightly below the prediction of Li et al. (2023) with maximum star formation efficiency = 0.2, and significantly below the upper limit of = 1.
@@ 6.5 Stellar Mass Density Evolution in the Early Universe
The growth of stellar mass density in the early Universe is highly uncertain. Some observational studies (e.g. Oesch et al., 2014; Stefanon et al., 2021; Willott et al., 2023a) have found a sharp decline in stellar mass density at 8, whereas others see a flatter evolution (Kikuchihara et al., 2020; Bhatawdekar et al., 2019). On the theoretical side, Oesch et al. (2018) uses dark matter halo evolutionary models to predict a deviation from the constant star formation efficiency (CSFE) model of Madau & Haardt (2015), which follows a significantly steeper slope at 7. Ferrara et al. (2023) has also recently presented a new model for the UV LF and SMF incorporating obscured star formation as well as a lack of dust within high- galaxies, which most closely matches our SMD results, despite predicting a steeper slope than we observe at .
@@
The range of stellar mass densities possible with our alternative GSMF estimates at is mostly above the 1 range of Adams et al. (2023). A discrepancy between the integrated star formation rate density and stellar mass density measured for the same sample could hint at a different IMF, since the assumed return fraction is strongly dependent on the chosen IMF, and the SMF and UVLF probe different stellar populations with different characteristic stellar mass. However there are a number of other possible issues with the conversion of the UVLF into a SMD estimate; the conversion of UV flux to SFR (), is often assumed to be a constant factor but is actually dependent on the age and metallicity Madau & Dickinson (2014), as well as the other assumptions used to calculate the return fraction (closed-box model, constant IMF & metal yield and instantaneous recycling of metals) which may not be valid at high-redshift.